A proton, moving with a velocity of 21, collides elastically with another proton that is initially rest. If the two protons have equal speeds, after the collision. Q7:- Find the speed of each proton after the collision (A)0.71ms-1 (B) 1.4ms (C) 2.1ms1 (D) 2.8ms1(E) 3.5ms-1 Q8:-The direction of the velocity vectors after the collision. (A)(SS,-35) (B)(30,-60) (C)(45,-45) (D)(60,-30) (E) (50,-40 )
Answer #1

7)

As collision is elastic, they conserve energy

1/2 mu^2 = 1/2 x 2m x v^2

u^2 = 2v^2

2^2 = 2xv^2

v = 1.4 m/s

8)

C